OpenStudy (anonymous):

SImplify: ³/16 + ³/54 (Third root of 16 + third root of 54)

OpenStudy (anonymous):

\[\sqrt[3]{16}+\sqrt[3]{54}\]

OpenStudy (amistre64):

3rt(16) + 3rt(54) might be easier to write :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (amistre64):

to add them you need to get like terms under the radicals

OpenStudy (anonymous):

how do I do that?

OpenStudy (amistre64):

like factors if you will

OpenStudy (amistre64):

hmmm, try converting to exponents and see if we can play with those maybe?

OpenStudy (amistre64):

\[16^{1/3}+54^{1/3}\] hmm

OpenStudy (anonymous):

so like: \[16^{1/3} + 54^{1/3}\]

OpenStudy (anonymous):

= 3rt (8*2) + 3rt(27*2) = 2 3rt2 + 3 3rt2 = 5 3rt2

OpenStudy (amistre64):

4*4 = 16 that aint it 9*6 = 54 .... thatll do it

OpenStudy (amistre64):

not mine; jimmys lol

OpenStudy (amistre64):

might help if I could recall the perfect cubes :/

OpenStudy (anonymous):

THanks jimmy's was right, but I don't know how he too hte 2 and the 3 from under the radical sign???

OpenStudy (anonymous):

how he took the**

OpenStudy (anonymous):

oh! never mind thanks :)